OpenStudy (anonymous):

Could someone help verify these answers i ahve: i was asked to find the abso/rel. max and min for this function: y=x^3-3x^2+4 over the interval [-2,4]. I have the rel. max as being 4, the rel. min as 0. Then i have the abs.min as -18 and abso. max as 20.

OpenStudy (anonymous):

Am i correct, with what i have

OpenStudy (anonymous):

Are you sure what you have is correct?

OpenStudy (anonymous):

on your first post x shoudl equal 2 not 4

myininaya (myininaya):

yes im wrong

myininaya (myininaya):

y'=3x(x-2) right lol i said x=4 not x=4 that was an oops x=0 x=2

myininaya (myininaya):

x=0, y=4 x=2, y=2^3-3(2^2)+4=8-12+4=0

myininaya (myininaya):

y''=6x-6 y''(0)=-6 => max y''(2)=6(2)-6=12-6=6 => min

myininaya (myininaya):

you can also use the first deriavtive y'=3x(x-2) and use test points around the critical numbers intead of using 2nd derivative

myininaya (myininaya):

to find max and mins

OpenStudy (anonymous):

Rel. max: 4 at x=0, Rel. min: 0 at x=2 Abso max: 20 at x=4, Abso min: -18 at -2 Is this correct?

myininaya (myininaya):

f'(-1)=+ f'(1)=- f'(3)=+ ---------|------|---- 0 2 ^ U so max at x=0 and min at x=2

OpenStudy (anonymous):

now is that the relative max and min or abso max and min

OpenStudy (anonymous):

or is it both, since 0 and 2 occur within the invtervals

myininaya (myininaya):

oh we are looking at an interval

myininaya (myininaya):

x=0, y=4 x=2, y=0 x=-2, y=-16 abs min at x=-2 x=4, y=20 abs max at x=4

OpenStudy (anonymous):

so is there no relative min and max at 2 and 0

myininaya (myininaya):

there is i already said they were above did you already forget lol?

myininaya (myininaya):

i said rel min at x=2 and rel max at x=0

OpenStudy (anonymous):

no, i didnt forget, i just want ot make sure that there is an indeed a rel max and min, and abso max and min,

OpenStudy (anonymous):

your funny

myininaya (myininaya):

no u lol

myininaya (myininaya):

so do you have any questions?

myininaya (myininaya):

i didn't read your whole paragraph above at first so i didn't know we were just looking at an interval

OpenStudy (anonymous):

find max and min of y=(x-1)^2(x+1)^2 smarty pants